# Deep Learning for Pathological Speech: A Survey

### v1

@@ IV Automatic Pathological Speech Detection
ParkCeleb 40 40 PD English Speech (L) Accessible
@@
Neurodegenerative diseases such as PD and ALS can significantly impact various sensorimotor and articulatory processes, leading to pathological speech impairments such as dysarthria, apraxia of speech, aphasia, dysphonia. As discussed in Section I, clinicians traditionally rely on laborious and time-consuming auditory-perceptual measures to diagnose these speech impairments accurately. To address the various challenges associated with auditory-perceptual assessments, there has been a growing interest in the research community to develop automatic approaches for diagnosing pathological speech. The goal of these approaches is to enhance the detection accuracy to match or surpass human accuracy, thereby ensuring more consistent and reliable diagnoses. Automated systems leverage advanced technologies such as classical machine learning and deep learning models with input representations as discussed in Section II to analyze speech patterns and identify anomalies indicative of these neurodegenerative conditions. In the following, we briefly summarize various classical machine learning-based and deep learning-based approaches. Additionally, insights on the speech modes used for pathological speech detection are provided.
@@ IV-A Classical Machine Learning-based Approaches
[126, 127] employed random forests and support vector machines (SVM), reporting promising results on a small pool of speakers. Similarly, [128] demonstrated good performance in detecting pathological dysarthric speech in an Indian Tamil language dataset using Gaussian mixture model and hidden Markov model classifiers. [129] found that the use of articulatory features on a small set of 24 Czech speakers showed very good performance, but might not generalize well to larger sets of speakers. Despite the promising results from these methods using various spectral/cepstral features, it remains unclear which features are most discriminative in capturing pathological cues. [130] showed that the cepstral coefficients are more robust and achieved the highest performance on a Spanish dataset when used as input to the SVMs for detection of pathological speech. [131] exploits $63736373$ acoustic, $36003600$ articulatory, and $44$ sensory features for the detection of ALS using deep neural networks and showed state-of-the-art detection performance on a dataset consisting of $1111$ patients and $1111$ healthy speakers. In a similar fashion, [132] employed SVM with openSMILE features on a dataset of 123 Hebrew speakers. Additionally, [133] utilized glottal features and demonstrated that combining glottal features with openSMILE features enhances the performance of dysarthric speech detection. [45] introduced spectro-temporal sparsity features combined with SVMs, showing these features to be more effective discriminators for dysarthric speech compared to temporal sparsity features. [134] applied Grassmann discriminant analysis on spectro-temporal subspaces, using singular value decomposition based on clinical evidence of spectro-temporal distortions. Their models were evaluated on datasets-PC-GITA and MoSpeeDi.
The major limitation of such works is their focus on small speaker sets and bias towards factors such as gender, recording conditions, age, class, or language, leading to poor generalization in clinical scenarios. To investigate the impact of the distribution shift, [135] employed cross-database models for dysarthric pathological speech detection. They demonstrated that an SVM model trained on the UA-Speech dataset showed a significant performance drop when tested on the AMSDC dataset. This highlights the need for future work to develop models robust enough to handle distribution shifts, including variations in recording conditions and other factors. Furthermore, classical machine learning-based methods primarily rely on hand-crafted features, which may not effectively capture all abstract pathological cues. Finally, these methods also tend to overlook valuable metadata such as speaker identity, gender, and language.
@@ IV-B Deep Learning-based Approaches
[136, 30, 137] utilized convolutional neural networks (CNN) for dysarthric speech detection on a dataset consisting of 26 speakers, including 13 patients diagnosed with ALS and their gender-matched counterparts respectively. Although this method demonstrated promising results, there is significant variation in pathological detection at the individual speaker level. [46] employed phase spectra such as modified group delay and instantaneous frequency spectra, demonstrating that using these phase spectra as complementary features to magnitude spectra improves the performance of CNN models in detecting dysarthric speech. Similarly, [56] employed temporal envelope features with CNN for dysarthric classification. [138] used SVM and deep neural networks for the detection of pathological speech on a dataset of speakers diagnosed with ALS. While most proposed approaches are single-task approaches aiming at achieving only pathological speech detection, [139] introduced the first ever multi-task system where they trained jointly pathological classification (primary task) and other 11 articulatory deficit tasks. The representations learned during joint training are constrained by various speech tasks and shows better generalization in pathological speech detection across speakers on a Spanish dataset.
Most studies to date have not exploited the temporal nature of audio frames. [140, 51, 141] addressed this issue by introducing a long short-term memory (LSTM) temporal network for dysarthric speech classification. [140] evaluated LSTM on a Mandarin dataset comprising of 69 speakers, 31 diagnosed with dysarthria and the remaining as normal speakers, while [51] evaluated an LSTM-based model on the TORGO dataset. Although the methodology in [140] showed promising results, the model has been evaluated only on the syllables making it harder to generalize across spontaneous speech.
Despite the promising results achieved by the proposed methods using handcrafted acoustic features and time-frequency representations, their detection capabilities still fall short of human-level performance and remain inadequate for clinical applications. Instead of handcrafted acoustic features or time-frequency representations, researchers endeavored to leverage raw speech data and applied self-supervised learning models like w2v2 to derive meaningful contextual representations, with the goal of improving the detection performance of pathological speech [60, 142, 43]. To the best of our knowledge, such models have become the state-of-the-art models in pathological speech detection, outperforming other automatic approaches based on handcrafted acoustic features or time-frequency input representations. However, although advantageous, these models currently lack interpretability insights that might be useful for a clinician. Moreover, the models developed thus far are largely language-dependent, as the linguistic context plays a critical role in pathological speech. Developing language-agnostic pathological detection methods would be highly beneficial, allowing for universal application across patients worldwide [143].
IV-C Impact of Speech Mode on Pathological Speech Detection
The large majority of the previously reviewed automatic pathological speech detection approaches have been proposed and validated on controlled speech tasks.
Controlled speech, also known as non-spontaneous speech, involves utterances produced within a structured context. In this mode, neurotypical speakers and patients are given a phonetically balanced, carefully crafted text and asked to repeat it. The goal is to elicit pathological biomarkers and cues through these controlled utterances, which can offer insights into motor planning and execution. Tasks in this mode may include reading from a text and repeating specific words.
In contrast to controlled speech, non-controlled or spontaneous speech consists of unplanned utterances that mirror real-world communication. This type of speech requires greater cognitive effort for planning, motor execution, and articulation, making it more likely to reveal natural pathological biomarkers. Examples include storytelling, recounting incidents, or engaging in everyday conversations. The reliance on controlled speech, with its structured and phonetically balanced content, may not always capture the full spectrum of pathological cues. Incorporating spontaneous speech could provide a more comprehensive understanding of pathological speech patterns. It seems plausible that developing pathological speech detection methods for spontaneous speech could be highly beneficial, given its ease of collection [61]. Research in this area could make such methods more practical and effective, as spontaneous speech naturally reflects real-world communication and may offer a richer set of pathological cues.
@@ V Pathological Speech Recognition
Automatic speech recognition (ASR) systems have significantly advanced over the years, demonstrating impressive results in converting raw speech signals into their corresponding textual form. This has resulted in the widespread use of various interactive devices, including smartphones and voice assistants. However, they often fail to recognize low resource pathological speech, including speech from patients suffering from various neurodegenerative conditions such as PD or ALS [22].
@@
To alleviate this problem, several attempts have been made in improving the performance of ASR systems for low resource pathological speech. For example, [144] proposed a two-step speaker adaptation method. Initially, a model trained on extensive neurotypical speech data is fine-tuned for dysarthric ASR. Subsequently, this fine-tuned model undergoes further adaptation to the specific dysarthric speaker. Additionally, [145] proposed the additive angular margin loss to address intra-class variation among dysarthric speakers, demonstrating promising results with Japanese speakers.  [146] also showed that personalized ASR systems fine-tuned on pathological speech exhibit better recognition performance compared to speaker-independent ASR systems. However, the reliance on specific speakers in personalized ASR systems poses challenges for generalization settings.  [147] utilize lattice-free maximum mutual information to mitigate insertion errors, which are otherwise prevalent due to the slow speaking rates of individuals with dysarthria. Due to the limited amount of available pathological speech data, the progress of ASR for pathological speech remains restricted. In a different approach, [148] utilized transfer learning and found that speaker-based data selection leads to negative transfer. They recommended using utterance-based data selection with an entropy distribution to enhance recognition.  [149] further implemented a multi-task learning (MTL) framework through auto-encoder joint learning, utilizing bottleneck features on out-of-domain data. Employing MTL in pathological ASR showed a lower word error rate compared to its single-task counterpart.  [150] demonstrated that state-of-the-art ASR systems for pathological speech are significantly impacted by phoneme inaccuracies. To address this, they proposed Speech Vision—a transfer learning paradigm that converts word utterances into visual feature representations, aiming to recognize the shape of the word rather than relying on phonemes. A comprehensive overview of ASR systems for pathological speakers in terms of progress and challenges is provided in [23].
@@
Due to the large number of parameters in state-of-the-art ASR models, fine-tuning them on low-resource pathological speech data is highly expensive. However, leveraging fine-tuning techniques such as LoRA (Low-Rank Adaptation) [151, 152] offers a more efficient alternative, reducing computational costs and making training more feasible. Additionally, LoRA’s modular approach might possibly make it easier to adapt the fine-tuned models to new speakers, providing flexibility in real-world applications or clinical scenarios.

### v2

@@ II Pathological Speech Datasets
    EWA-DB 896 226 PD, AD, MCI Slovak Speech Accessible
    ParkCeleb 40 40 PD English Speech + Video + L Accessible
    mPower 5,581 1,087 PD English Multimodal Accessible
  • PC-GITA [53].  The PC-GITA dataset contains Spanish (spontaneous and read) speech recordings from $50$ control speakers and $50$ patients suffering from PD. The two groups of speakers are age- and sex-matched, with $25$ male and $25$ female speakers in each group.
  • MoSpeeDi [54].  The MoSpeeDi dataset contains French (spontaneous and read) speech recordings from $466$ control speakers and $138$ patients suffering from various types of motor speech disorders such as dysarthria or apraxia of speech. While subgroups of age- and sex-matched controls and patients can be found within the database, the overall dataset is not age- and sex-matched.
  • Nemours [55].  The Nemours dataset is an English dataset containing (read) speech recordings from $11$ male speakers with varying degrees of dysarthria severity.
  • CUDYS [56].  The Cantonese Dysarthric Speech Corpus contains Cantonese (read) speech recordings from $5$ control speakers and $11$ patients suffering from cerebellar degeneration. The control group consists of $2$ female and $3$ male speakers, whereas the dysarthric group consists of $5$ female and $6$ male speakers.
  • AMSDC [57].  The Atlanta Motor Speech Disorders Corpus is an English dataset containing (spontaneous and read) speech recordings from $99$ ($62$ male, $37$ female) patients diagnosed with various disorders such as PD, ALS, or dementia.
  • EST [58].  The EST dataset is a Dutch dataset containing (read) speech recordings from $16$ male dysarthric patients due to PD, traumatic brain injuries, or cerebrovascular accident. .
  • EasyCall [59].  The EasyCall dataset is an Italian dataset containing (read) speech recordings from $24$ controls and $31$ patients diagnosed with PD, Huntington’s disease, ALS, and peripheral neuropathy. The control group consists of $10$ female and $14$ male speakers, whereas the patient group consists of $11$ female and $20$ male speakers.
  • COPAS [60].  The Corpus of Pathological and Normal Speech dataset is a Dutch dataset containing (spontaneous and read) speech recordings from $197$ pathological speakers and $122$ control speakers. The database comprises $8$ distinct pathological categories such as dysarthria, hearing impairment, cleft, etc.
  • ParkCeleb [61].  The previously reviewed datasets are not longitudinal and do not allow tracking the progression of the speech disorder within the same patient along time. To address this gap, the English ParkCeleb dataset was recently introduced in [61]. This dataset contains (spontaneous) audio-visual recordings (such as studio interviews or press conferences) from $40$ (22 female, $38$ male) control speakers and $40$ (22 female, $38$ male) patients suffering from PD.
  • Italian Parkinson’s Database [62]. The dataset contains Italian (read) speech recordings from a total of $65$ speakers, including $28$ patients with PD and $37$ control speakers. Sex distribution is non-uniform across groups, with the PD group consisting of $19$ male and $9$ female speakers and the control group consisting of $23$ male and $14$ female speakers.
  • NeuroVoz [63]. The NeuroVoz dataset contains ( spontaneous, read, and listen and repeat) speech recordings from $112$ Castilian Spanish speakers, including $54$ patients diagnosed with PD and $58$ control speakers. The control group consists of $28$ male speakers, $26$ female speakers, and $1$ speaker whose sex information is not provided, whereas the patient group consists of $33$ male and $20$ female speakers.
  • Saarbrücken Voice Database [64] [^2] https://stimmdb.coli.uni-saarland.de/ (accessed July 07, 2025). The Saarbrücken Voice Database contains (read) speech recordings and electroglottography data from $2,255$ German speakers, including $1,356$ patients and $869$ control speakers [65]. Patients are diagnosed with various voice disorders. The control group consists of $433$ male and $436$ female speakers, whereas the patient group consists of $629$ male and $727$ female speakers.
  • Turkish Parkinson Speech Dataset [66]. This dataset contains (read) speech recordings from $40$ Turkish speakers, including $20$ controls and $20$ patients with PD. The PD group contains $6$ female and $14$ male speakers, whereas the control group contains $10$ female and $10$ male speakers.
  • Czech Parkinson’s Dataset [67]. This dataset contains (read) speech recordings from $83$ Czech speakers, including $22$ control speakers and $61$ patients diagnosed with PD or atypical parkinsonian syndromes. The patient group consists of $30$ female and $31$ male speakers, whereas the control group consists of $11$ female and $11$ male speakers.
  • EWA-DB [68]. The EWA-DB dataset consists of (spontaneous and read) speech recordings of $1,122$ Slovak speakers, including $896$ control speakers and $226$ patients diagnosed with PD, mild cognitive impairment, or Alzheimer’s disease. The patient group consists of $121$ male and $105$ female speakers, whereas the control group consists of $248$ male and $648$ female speakers.
  • mPower Parkinson’s Dataset [69]. The mPower study collected multimodal smartphone sensor data from $6805$ participants, including $1,087$ patients diagnosed with PD and $5,581$ control speakers. Patient and control status were self-reported. The dataset was collected through the mPower mobile application and includes four distinct sensor-based assessment modalities, i.e., spatial memory evaluation, gait analysis through walking tasks, manual dexterity measurement via finger tapping, and vocal function assessment using sustained phonation recordings.
As presented in Table I, a considerable number of the datasets used in the literature are private and not openly accessible. This limits their availability for broader research and replication efforts. Although some datasets such as TORGO are publicly available, they suffer from detrimental recording artifacts [70], which compromise their usefulness for studying pathological speech characteristics and for developing reliable automatic detection methods [70, 71]. The lack of accessible, high-quality data poses a major challenge to the development and validation of effective algorithms, slowing progress in understanding and addressing pathological speech conditions. Ensuring the availability of clean, reliable, and comprehensive datasets is therefore essential for advancing research and practical applications.
Another critical limitation is the lack of linguistic diversity. Existing datasets are available only in a few languages, which restricts their applicability to non-represented populations and limits the potential for cross-linguistic comparisons. This linguistic bias poses a challenge for developing universal models that can generalize across languages.
In addition to these issues, most datasets are relatively small, often including only a limited number of participants. This scarcity hinders the development of robust and generalizable models and makes it difficult to capture the wide variability in speech disorders across different individuals and conditions. Furthermore, demographic imbalances are common. Many datasets exhibit skewed gender representation, which may introduce biases into trained models. In some cases, age groups or types of pathological conditions are also underrepresented, further limiting generalizability.
These limitations underscore the pressing need for larger, more diverse, and better-balanced pathological speech datasets that reflect the complexity of real-world populations. Addressing these gaps is essential for building reliable and inclusive tools for pathological speech analysis.
III Speech Representations for
Automatic Approaches
Pathological speech exhibits a range of acoustic anomalies, including deviations in pitch, loudness, vowel space reduction, and articulation [72, 4]. Additionally, it can lead to asymmetrical tension in the vocal folds, resulting in irregular vibrations and, consequently, an abnormal fundamental frequency. Patients can also show inconsistent rhythmic structures in comparison to control groups [73]. Excessively high or low fundamental frequency, combined with excessive vocal intensity, can exacerbate the severity of pathological voice conditions, producing characteristics such as shrillness, screechiness, hoarseness, or huskiness [74].
To capture such abnormalities as biomarkers for automatic pathological speech processing, researchers have employed various handcrafted acoustic features such as OpenSMILE [70], Mel-frequency cepstral coefficients (MFCCs) [75], or spectro-temporal sparsity features [76]. Raw speech signals and various time-frequency representations such as the short-time Fourier transform (STFT) have also been directly exploited in combination with deep learning approaches to directly learn pathology-discriminant cues [77, 78]. More recently, latent embeddings derived from self-supervised models have been used as more powerful representations of speech patterns and the various impairments [70, 79, 80]. It is important to note that a considerably larger set of features and representations have been employed to characterize disorders across various pathologies than the ones outlined above. Here, we briefly discuss the representations that are most commonly used in the literature.
III-A Handcrafted Acoustic Features
OpenSMILE
Among the various features used in pathological speech detection, OpenSMILE features have been widely explored in the literature in conjunction with traditional machine learning algorithms [81, 82, 83, 84, 40, 85, 86, 87, 88, 89, 70, 90, 91, 92]. The OpenSMILE feature set includes a $6552$-dimensional feature vector that primarily consists of low-level audio features such as CHROMA, CENS, loudness, MFCCs, and other spectral features. In the context of pathological speech, these features can capture the subtle anomalies in speech patterns that are indicative of disorders. OpenSMILE’s comprehensive feature set allows for the analysis of prosodic elements like pitch, jitter, shimmer, and formant frequencies, which are crucial for identifying and differentiating various speech pathologies. However, OpenSMILE features are general features that have been used for a variety of speech applications such as emotion recognition [93] or speech recognition [94] and they are not specifically handcrafted to capture pathological cues..
Spectral and cepstral coefficients
Due to their ability to characterize articulation deficiencies, various spectral and cepstral coefficients such as Linear Predictive Coding (LPC), MFCCs, and Perceptual Linear Prediction (PLP) features, have been successfully used for pathological speech detection [95, 96, 97, 98, 99, 100, 101, 102, 103, 13, 104]. LPC features describe the distribution of energy across frequency bands and are directly related to the resonant properties of the vocal tract, making them suitable for analyzing articulatory features like formant frequencies and tongue positioning. In contrast, MFCCs and PLP features are obtained by transforming the spectral envelope into the cepstral domain, which effectively separates source and filter components while capturing smoothed representations of the vocal tract shape. When extended with temporal dynamics, these features can track transitions and instability in articulation patterns. To enhance robustness against channel variability and noise, the Relative Spectral Transform PLP features have also been considered in [105, 106]. Furthermore, the fusion of modulation spectra features or glottal features with MFCCs has additionally been explored to further improve performance [107, 108].
Spectro-temporal sparsity
III-B Time-Frequency Representations
III-C Raw Waveform Representation
While handcrafted acoustic features and time-frequency representations have demonstrated strong performance, researchers have also explored end-to-end pathological speech detection using raw input representations, eliminating the need for handcrafted features or time-frequency representations. Studies such as [78, 123, 124] have shown that raw waveform-based methods can capture discriminative pathological patterns and outperform traditional feature-based approaches. However, these methods may require more training data compared to other approaches for a robust performance.
III-D Self Supervised Embeddings
Even though handcrafted acoustic features, time-frequency, and raw waveform input representations have achieved promising results in the analysis of pathological speech, their performance remains limited. The current state-of-the-art input representations for pathological speech primarily consist of latent embeddings derived from self-supervised models such as wav2vec2 [125], HuBERT [126], or WaveLM [127]. Research in e.g., [128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 79, 141, 142, 143, 144, 92, 145, 146, 147, 148, 149, 150, 151] has shown that these advanced models are effective in capturing complex speech patterns and nuances, leading to improved performance in tasks like detection and recognition. However, ongoing research continues to refine these embeddings and assess their robustness and generalizability across diverse pathological speech conditions, tasks, and applications.
IV Automatic Pathological Speech Detection
@@
As discussed in Section I, clinicians traditionally rely on laborious and time-consuming auditory-perceptual measures to diagnose speech impairments accurately. To address the various challenges associated with auditory-perceptual assessments, there has been a growing interest in the research community to develop automatic approaches for diagnosing pathological speech. The goal of these approaches is to enhance the detection accuracy to match or surpass human accuracy, thereby ensuring more consistent and reliable diagnosis. Automated systems analyze speech patterns and identify anomalies indicative of these neurodegenerative conditions by leveraging classical machine learning and deep learning models with input representations such as the ones discussed in Section III. In the following, we briefly summarize various classical machine learning-based and deep learning-based approaches.
@@ IV-A Classical Machine Learning-based Approaches
In recent years, numerous studies have investigated the use of classical machine learning classifiers combined with hand-crafted acoustic features for detecting pathological speech, yielding promising results. For instance, [152] and [153] employed Random Forests and Support Vector Machines (SVMs) on a range of dysphonia measures, achieving strong classification performance on relatively small speaker datasets. Similarly, [154] applied Gaussian Mixture Models and Hidden Markov Models to cepstral features for detecting dysarthric speech in an Indian Tamil language dataset. Orozco‐Arroyave et al. [155] further demonstrated that cepstral coefficients are effective and robust when used with SVMs for detecting dysarthric speech in a Spanish dataset. While these findings are encouraging, their scalability to larger, more diverse datasets remains uncertain due to the variability of speech pathologies across populations. Additionally, despite the effectiveness of various spectral and cepstral features, the field still lacks consensus on which features are the most discriminative and generalizable across different conditions and datasets.
Recognizing the importance of developing discriminative feature representations, the research community has devoted significant efforts to this area. For example, [156] investigated the use of articulatory features for pathological speech detection and achieved promising results on a small dataset of 24 Czech speakers. However, the limited sample size restricts the generalizability of the findings, and extracting articulatory features at scale remains technically challenging. In a more comprehensive effort, [157] utilized an extensive feature set including $6,373$ acoustic, $3,600$ articulatory, and $4$ sensory features to detect ALS, reporting state-of-the-art performance on a dataset of $22$ speakers. Norel et al. [158] employed SVMs with openSMILE features to detect pathological speech in a larger cohort of 123 Hebrew speakers. Prabhakera and Alku [82] found that incorporating glottal features alongside openSMILE features improved dysarthric speech detection performance. Further advancements were made in [76] by introducing spectro-temporal sparsity features, which outperformed temporal sparsity features in classifying dysarthric speech using SVMs. Additionally, [159] applied Grassmann discriminant analysis to spectro-temporal subspaces, leveraging singular value decomposition informed by clinical insights into pathological distortions.
A common limitation across these studies is their reliance on relatively small speaker sets and their susceptibility to biases related to sex, age, recording conditions, or language. To evaluate generalization across different domains, [160] conducted a cross-database study, demonstrating that an SVM trained on one dataset suffered a significant performance drop when evaluated on another dataset. This highlights the critical need for models that are robust to distributional shifts, including those stemming from differences in dataset characteristics and recording environments. Moreover, classical machine learning approaches heavily depend on hand-crafted features, which may not fully capture the nuanced and abstract cues associated with pathological speech. These methods also tend to overlook important metadata such as speaker identity, sex, and language.
@@ IV-B Deep Learning-based Approaches
Several studies have explored CNN-based architectures for pathological speech detection, leveraging their ability to extract local patterns from time-frequency representations. For instance, [39, 161] applied CNNs to dysarthric speech, achieving promising results on small datasets of control speakers and ALS patients. However, these models exhibited significant inter-speaker variability in performance, suggesting a lack of robustness across individuals. This limitation is particularly critical in clinical contexts where model reliability must generalize across diverse patient populations. Building on CNN-based models, [77] incorporated phase-based features, i.e., the modified group delay and instantaneous frequency spectra, as complementary inputs to magnitude spectra. This multi-representational approach improved the CNN performance for pathological speech detection. Similarly, [88] introduced temporal envelope features, reinforcing that temporal dynamics encode critical pathological cues. However, these representation augmentations increase model complexity and demand careful preprocessing, which may hinder real-world deployment. While most approaches focus solely on binary classification of pathological versus neurotypical speech, multi-task learning has shown potential for capturing broader articulatory patterns. Vásquez Correa et al. [162] introduced a multi-task system based on CNNs that simultaneously learned to classify pathological speech and predict $11$ articulatory deficit attributes. This approach yielded improved generalization across speakers by constraining the learned representations with related speech tasks, demonstrating the utility of auxiliary objectives in enhancing model robustness.
While CNNs excel at spatial pattern recognition, they lack mechanisms to capture temporal dependencies inherent in speech. Recurrent architectures, particularly LSTMs, have been employed to address this gap [163, 83, 164, 123]. These models demonstrate improved performance but often require larger datasets to train effectively. Moreover, although the LSTM methodology in [163] showed promising results, the model has been evaluated only on syllables and its generalization to other speech modes remains unexplored.
More recently, there has been a shift toward end-to-end models using self-supervised learning (SSL) to bypass manual feature engineering. SSL models like wav2vec 2.0 leverage raw waveform inputs to learn contextual embeddings, achieving state-of-the-art results across several pathological speech datasets [91, 165, 70]. These models outperform classical approaches in both accuracy and scalability. However, despite their empirical success, SSL-based models function as black boxes, raising concerns about clinical interpretability and decision transparency.
@@ V Pathological Speech Recognition
ASR systems have significantly advanced over the years, demonstrating impressive results in converting raw speech signals into their corresponding textual form. This has resulted in the widespread use of various interactive devices, including smartphones and voice assistants. However, they often fail to recognize low resource pathological speech, including speech from patients suffering from various neurodegenerative conditions such as PD or ALS [31].
@@
To alleviate this problem, several attempts have been made in improving the performance of ASR systems for pathological speakers. For example, [166] proposed a two-step speaker adaptation method. In the first step, a model trained on extensive control speech data is fine-tuned for dysarthric ASR. In the second step, this fine-tuned model undergoes further adaptation to a specific dysarthric speaker. Additionally, [167] proposed the additive angular margin loss to address intra-class variation among dysarthric speakers, demonstrating promising results on Japanese speakers.  Green et al. [168] also showed that personalized ASR systems fine-tuned on pathological speech exhibit better recognition performance compared to speaker-independent ASR systems. However, the reliance on specific speakers in personalized ASR systems poses challenges for generalization settings. Hermann and Magimai.-Doss [169] utilize lattice-free maximum mutual information to mitigate insertion errors, which are otherwise prevalent due to the slow speaking rates of individuals with dysarthria. In a different approach, Xiong et al. [170] utilized transfer learning and found that speaker-based data selection leads to negative transfer. They recommended using utterance-based data selection with an entropy distribution to enhance recognition. Yue et al. [171] further implemented a multi-task learning (MTL) framework through auto-encoder joint learning, utilizing bottleneck features on out-of-domain data. Employing MTL in pathological ASR showed a lower word error rate compared to its single-task counterpart.  Shahamiri [172] demonstrated that state-of-the-art ASR systems for pathological speech are significantly impacted by phoneme inaccuracies. To address this, they proposed Speech Vision, a transfer learning paradigm that converts word utterances into visual feature representations, aiming to recognize the shape of the word rather than relying on phonemes.
@@
A comprehensive overview of ASR systems for pathological speakers in terms of progress and challenges is provided in [32], where it is shown that pathological speech recognition performance can be significantly improved through a combination of neural architecture search, data augmentation, speaker adaptation, and multi-modal learning. These techniques address challenges such as limited training data, high inter-speaker variability, and reduced speech intelligibility. It should be noted that due to the large number of parameters in state-of-the-art ASR models, fine-tuning them on low-resource pathological speech data can be highly expensive. However, leveraging fine-tuning techniques such as Low-Rank Adaptation (LoRA) offers a more efficient alternative, reducing computational costs and making training more feasible [173, 174]. Additionally, LoRA’s modular approach might make it easier to adapt the fine-tuned models to new speakers, providing flexibility in real-world applications or clinical scenarios.